OpenStudy (anonymous):

if U = 3i-4j, and V = 2i+3j then the angle Ө between U and V are ?

OpenStudy (anonymous):

Have you guys learned dot or cross products in your class?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

you can pick either one. i prefer dot products though

OpenStudy (anonymous):

so it's U dot V = ac + bd ?

OpenStudy (anonymous):

U dot V = 3(2) + -4(3) ?

OpenStudy (anonymous):

correct. U dot V = ac + bd = |u||v|cos(theta)

OpenStudy (anonymous):

are those 2 formulas the same ?

OpenStudy (anonymous):

yes, they are equivalent. you can just write it as ac+bd = |u||v|cos(theta)

OpenStudy (anonymous):

solve for theta and you got your angle :)

OpenStudy (anonymous):

I'm sorry i got -6 for U dot V then what do i do from there ?

OpenStudy (anonymous):

|u| is the magnitude of the u vector. so 3^2 + 4 ^ 2 = c ^2, solve for c and get 5. |v| is the magnitude of the v vector. so 2^2 + 3^2 = c^2 solve for c and get sqrt(13)

OpenStudy (anonymous):

-6 = 5* sqrt(13) * cos(theta)

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

so the answer is : cos(theta) = 11xsqrt(13) that's it ?

OpenStudy (anonymous):

\[\cos(\theta)=-6/(-5*\sqrt(13))\] \[\theta = \cos^{-1} (-6/(5*\sqrt(13))\]

OpenStudy (anonymous):

i don't have a calculator on me :( but that should give you your answer

OpenStudy (anonymous):

how did u move that -6 o_0 and the cos theta

OpenStudy (anonymous):

don't you add the -6 to the right side ? and minus cos(theta) to the left side ?

OpenStudy (anonymous):

ahh nvm i got it, you used all 3 formulas at the sametime, and just plugging it in =.=

OpenStudy (anonymous):

Thanks @alexandercpark